OpenStudy (anonymous):

use the multiplication property of inequalities to solve the inequality \[\frac{ 3 }{ 4 }x \ge -\frac{ 1 }{ 8 }\]...what is the solution set??

OpenStudy (amriju):

hey cud u use 'equation' to write the ques..cant understand the equation...

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }x \ge -\frac{ 1 }{ 8 }\] @amriju

OpenStudy (amriju):

sorry...a lot of \ ..ge...i cant understand...